Solve by using elimination:
x-y-2z=4
-x+2y+z=1
-x+y-3z=11
Please and thank you

Answer by Alan3354(69443)

Solve by using elimination:
x-y-2z=4
-x+2y+z=1
-x+y-3z=11
--------------
Eliminate x first.
---
Add eqn 1, to eqn 2.
Then eqn 1 to eqn 3.